calculate the Maximum amount of p-aminophenol/5ml permitted in the acetaminophen 100mg/5ml. USP limit for p-aminophenol is 0.005% w/v
Answer #1

USP limit = 0.005% w/v = 0.005 g / 100 mL

In 5 mL, the USP limit is : 0.005 / 20 = 5 * 10-3 / 20 = 2.5 * 10-4 g / 5 mL = 0.25 mg / 5 mL

Therefore the maximum amount permitted is 0.25 mg per 5 mL